OpenStudy (anonymous):

what two numbers whose product is 40 and whose sum is -14

OpenStudy (ajprincess):

x*y=40 x+y=-14 x=-14-y y(-14-y)=40 -14y-y^2=40 y^2+14y+40=0 Use quadratic formula to find y.